OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -12, -7, -2, ...

OpenStudy (anonymous):

an = -17 + 5(n + 2) an = -17 + 5(n + 1) an = -17 + 5(n - 1) an = -17 x 5(n - 1)

OpenStudy (anonymous):

those arent an option

OpenStudy (primeralph):